### v2

# Right on Time: Revising Time Series Models by Constraining their Explanations

Abstract
The reliability of deep time series models is often compromised by their tendency to rely on confounding factors, which may lead to misleading results. Our newly recorded, naturally confounded dataset named P2S from a real mechanical production line emphasizes this. To tackle the challenging problem of mitigating confounders in time series data, we introduce Right on Time (RioT). Our method enables interactions with model explanations across both the time and frequency domain. Feedback on explanations in both domains is then used to constrain the model, steering it away from the annotated confounding factors. The dual-domain interaction strategy is crucial for effectively addressing confounders in time series datasets. We empirically demonstrate that RioT can effectively guide models away from the wrong reasons in P2S as well as popular time series classification and forecasting datasets.
Machine Learning, Time Series, XAI
1 Introduction
Time series data is ubiquitous in our world today. Everything that is measured over time generates some form of time series, for example, energy load (Koprinska et al., 2018), sensor measurements in industrial machinery (Mehdiyev et al., 2017) or recordings of traffic data (Ma et al., 2022). Various neural models are often applied to complex time series data (Ruiz et al., 2021; Benidis et al., 2023). As in other domains, these can be subject to confounding factors, ranging from simple noise or artifacts to complex shortcut confounders (Lapuschkin et al., 2019). However, confounders and the mitigation there of have yet to receive attention for time series, where existing works make specific assumptions about settings and data for success (Bica et al., 2020).
In particular, the mitigation of shortcut confounders, i.e., spurious patterns in the training data used for the prediction, is essential. If a model utilizes confounding factors in the training set, its decision relies on wrong reasons, letting it fail to generalize to data without the confounder. Model explanations play a crucial role in uncovering confounding factors, but they are not enough on their own to address them. While an explanation can reveal that the model relies on incorrect factors, it does not alter the model’s outcome. To change this, we introduce Right on Time (RioT), a new method following the core ideas of explanatory interactive learning (XIL) (Teso & Kersting, 2019), i.e., using feedback on explanations to mitigate confounders. RioT uses traditional explanation methods like Integrated Gradients (IG) (Sundararajan et al., 2017) to detect whether models focus on the right or the wrong time steps and utilizes feedback on the latter to revise the model (Fig. 1, top).
Confounding factors in time series data are not limited to the time domain. A steady noise frequency in an audio signal can also be a confounder but cannot be pinned to a specific time step. To handle these kinds of confounders, RioT can also incorporate feedback in the frequency domain (Fig. 1, bottom). To further emphasize the importance of mitigating confounders in time series data, we introduce a new real-world, confounded dataset called P2S. The dataset consists of sensor measurements from an industrial high-speed press, which is part of many important manufacturing processes in the sheet metal working industry. The sensor data used to detect faulty production is naturally confounded and thus causes incorrect predictions after training. Next to its immediate industrial relevance, P2S is the first time series dataset that contains explicitly annotated confounders, enabling evaluation and comparison of confounder mitigation strategies on real data.
Figure 1: Top: Expert annotations marking the wrong reasons (red) are given based on model explanations to mitigate spatial confounders. The application of RioTsp revises the model, whose adjusted explanations reveal that its output is based on the right reasons (blue) instead. Bottom: As confounders can also occur in the frequency domain, RioTfreq can use explanations and feedback in this domain as well. Similar to the spatial domain, it guides the model’s focus, mitigating the influence of the confounders.
Altogether, we make the following contributions: (1) We show both on our newly introduced real-world dataset P2S and on several other manually confounded datasets that SOTA neural networks on time series classification and forecasting can be affected by confounders. (2) We introduce RioT to mitigate confounders for time series data. The method can incorporate feedback not only on the time domain but also on the frequency domain. (3) By incorporating explanations and feedback in the frequency domain, we enable a new perspective on XIL, overcoming the important limitation that confounders must be spatially separable. [^1] Code will be available at a later point in time
The remainder of the paper is structured as follows. In Sec. 2, a brief overview of related work on explanations for time series and revising model mistakes is given. In Sec. 3, we introduce our method before providing an in-depth evaluation and discussion of the results in Sec. 4. Lastly, in Sec. 5, we conclude the paper and provide some potential avenues for future research.
2 Related Work
Our work touches upon fields of explainable AI and time series prediction. Let us provide a brief overview of these.
Explanations for Time Series. Within the field of explainable artificial intelligence (XAI), various techniques to explain machine learning models and their outcomes have been proposed. While many techniques originate from image or text data, they were quickly adapted to time series (Rojat et al., 2021). While backpropagation- and perturbation-based attribution methods provide explanations directly in the input space, other techniques like symbol aggregations (Lin et al., 2003) or shapelets (Ye & Keogh, 2011) aim to provide higher-level explanations. For a more in-depth discussion of explanations for time series, we refer to surveys by Rojat et al. (2021) or Schlegel et al. (2019). While it is essential to have explanation methods to detect confounding factors, they alone are insufficient to revise a model. Thus, explanations are the starting point of our method, as they enable users to detect confounders and provide feedback to overcome confounders in the model. In particular, we build upon Integrated Gradients (Sundararajan et al., 2017). This method computes attributions for the input by utilizing model gradients. We selected it because of its several desirable properties like completeness or implementation invariance and its wide use, also for time series data (Mercier et al., 2022; Veerappa et al., 2022).
Explanatory Interactive Learning (XIL). While not prevalent for time series data, there has been some work on confounders and how to overcome them for other domains, primarily the image domain. Most notable, there is explanatory interactive learning, which describes the general process of revising a model’s decision process based on human feedback (Teso & Kersting, 2019; Schramowski et al., 2020). Within XIL, the model’s explanations are used to incorporate the feedback back to the model, thus revising its mistakes (Friedrich et al., 2023a). XIL can, for example, be used on models that show Clever-Hans-like behavior to prevent them from using confounding shortcuts in the data (Stammer et al., 2020). Several methods apply the idea of XIL to image data. For example, Right for the Right Reasons (RRR) (Ross et al., 2017) and Right for Better Reasons (RBR) (Shao et al., 2021) use human feedback as a penalty mask on model explanations. Instead of penalizing wrong reasons, HINT (Selvaraju et al., 2019) rewards the model for focusing on the correct part of the input. Furthermore, Friedrich et al. (2023b) investigate the use of multiple explanation methods simultaneously. Although various XIL methods are often employed to address confounders in image data, their application to time series data remains unexplored. To bridge this gap, we introduce RioT, a method that incorporates the core principles of XIL to the unique characteristics of time series data.
Unconfounding Time Series. Next to approaches from interactive learning, there is also some other work on unconfounding time series models. This line of work is generally based on causal analysis of the time series model and data. Time Series Deconfounder (Bica et al., 2020) estimates treatment under the presence of hidden confounders. Similarly, Flanders et al. (2011) propose a method to detect residual confounders in time series of observational studies. These methods have specific assumptions about data, models, and confounders. This restricts their general usage, particularly to mitigate model mistakes after they have been identified via explanations. While this work also aims to mitigate confounders in time series data, our approach differs vastly from these works and has no specific assumptions about data and confounder besides that it is visible in explanations.
3 Right on Time (RioT)
The core intuition of Right on Time (RioT) is to utilize human feedback to steer a model away from wrong reasons. It follows the general structure of XIL, which has four main steps Friedrich et al. (2023a). In Select, instances for feedback and following model revision are selected. Following previous XIL methods, we select all samples by default while not necessarily requiring feedback for all of them. Afterwards, Explain covers how model explanations are generated, before in Obtain, a human provides feedback on the selected instances. Lastly, in Revise, the feedback is integrated into the model to overcome the confounders. We build RioT along these steps and cover how each is realized in the following. But let us first establish some notation for the remainder of this paper.
Given is a dataset $(\mathcal{X},\mathcal{Y})$ and a model $f(\cdot)$ for time series classification or forecasting. The dataset consists of $D$ many pairs of $\bm{x}$ and $\bm{y}$. Thereby, $\bm{x}\in\mathcal{X}$ is a time series of length $T$, i.e., $\bm{x}\in\mathbb{R}^{T}$ and $\bm{y}\in\mathcal{Y}$ is the ground-truth output, a class label $\bm{y}\in\{1,\dots,K\}$ for $K$ class classification or the ground-truth forecast $\bm{y}\in\mathbb{R}^{W}$ of length $W$. For a datapoint $\bm{x}$, the model generates the output $\hat{\bm{y}}=f(\bm{x})$, where the dimensions of $\hat{\bm{y}}$ are the same as of $\bm{y}$ for both tasks.
Figure 2: This figure depicts the flow of explanation and revision, beginning with input data $\bm{x}$, through the model $f(\bm{x})$ to explanations $e(\bm{x})$, annotated feedback $a(\bm{x})$, and finally back to the model. IG provides the spatial model explanation, and is transformed with the FFT to obtain the frequency explanation. Expert user annotations can be applied in either or both domains. They are utilized by the right-reason loss ($\mathcal{L}_{RR}^{sp}$ and $\mathcal{L}_{RR}^{freq}$) to guide the model away from confounders in both the time and frequency domains.
3.1 Explain
Given a pair of input $\bm{x}$ and model output $\hat{\bm{y}}$ for time series classification, the explainer generates an explanation $e_{f}(\bm{x})\in\mathbb{R}^{T}$ in the form of attributions to explain $\hat{\bm{y}}$ w.r.t. $\bm{x}$. For an element of the input, a large attribution value means a large influence of that element on the output. In the remainder of the paper, explanations refer to the model $f$, but we drop $f$ from the notation to declutter it, resulting in $e(\bm{x})$. For RioT, we use IG as explainer. However, we make some adjustments to the base method to make it more suitable for time series and model revision.
Part of IG is a multiplication of the model gradient with the input itself, improving the explanation’s quality (Shrikumar et al., 2017). However, this multiplication makes some implicit assumptions about the input format. In particular, it assumes that there are no inputs with negative values. Otherwise, multiplying the attribution score with a negative input would flip the attribution’s sign, which is not desired. For images, this is unproblematic because they are always equal to or larger than zero. In time series, negative values can occur and normalization to make them all positive is not always suitable. To avoid this problem, we use only the input magnitude and not the input sign to compute the IG attributions, resulting in the following equation:
$e(\bm{x})=|\bm{x}-\bar{\bm{x}}|\cdot\int_{0}^{1}\frac{\partial f(\tilde{\bm{x}})}{\partial\tilde{\bm{x}}}\Biggr{|}_{\tilde{\bm{x}}=\bar{\bm{x}}+\alpha(\bm{x}-\bar{\bm{x}})}d\alpha$ (1)
Attributions for Forecasting. In a classification setting, attributions are generated by propagating gradients back from the model output (of its highest activated class) to the model inputs. However, there is often no single model output in time series forecasting. Instead, the model generates one output for each timestep of the forecasting window simultaneously. Naively, one could use these $W$ outputs and generate as many explanations $e^{\prime}_{1}(\bm{x}),\dots e^{\prime}_{W}(\bm{x})$. This number of explanations would, however, make it even harder for humans to interpret the results, as the size of the explanation increases with $W$(Miller, 2019). Therefore, we propose to aggregate the individual explanations by averaging:
$e(\bm{x})=\frac{1}{W}\sum_{i=1}^{W}e^{\prime}_{i}(\bm{x})$ (2)
Averaging attributions over the forecasting window provides a simple yet robust aggregation of the explanations. Other means of combining them, potentially even weighted based on distance of the forecast in the future are also imaginable. Overall, this allows attributions for time series classification and forecasting to be generated similarly.
Attributions in the Frequency Domain. Time series data is often also given in a frequency representation. In some situations, this format can be more intuitive for humans to understand than spatial representations. As a result, providing explanations in this domain becomes essential. Additionally, frequency explanations enable feedback in this domain as well. Vielhaben et al. (2023) showed that it is possible to obtain frequency attributions of the method Layerwise Relevance Propagation (Bach et al., 2015), even if the model does not operate directly on the frequency domain. We transfer this idea to IG: for an input sample $\bm{x}$, we generate attributions with IG, resulting in $e(\bm{x})\in\mathbb{R}^{T}$. We then interpret the explanation as a time series, with the attribution scores as values. To obtain the frequency explanation, we perform a Fourier transformation of $e(\bm{x})$, resulting in the frequency explanation $\hat{e}(\bm{x})\in\mathbb{C}^{T}$ with $\hat{E}$ for the entire set.
3.2 Obtain
The next step of RioT is to obtain user feedback on confounding factors. For an input $\bm{x}$, a user can mark parts that are confounded, resulting in a feedback mask $a(\bm{x})\in\{0,1\}^{T}$. In this binary mask, a $1$ signals a potential confounder at this time step. Thereby, it is not necessary to have feedback for every sample of the dataset, as a mask $a(\bm{x})=(0,\dots,0)^{T}$ corresponds to no feedback. Feedback can also be given on the frequency explanation in a similar manner, marking which elements in the frequency domain are potential confounders. The resulting feedback mask $\hat{a}(\bm{x})=(\hat{a}(\bm{x})_{re},\hat{a}(\bm{x})_{im})$ can be different for the real $\hat{a}(\bm{x})_{re}\in\{0,1\}^{T}$ and imaginary part $\hat{a}(\bm{x})_{im}\in\{0,1\}^{T}$. For the whole dataset, we then have spatial annotations $A$ and frequency annotations $\hat{A}$. As the annotated feedback masks have to come from human experts, obtaining them can become costly. In many cases, however, it is possible to generate masks based on confounders for a couple of samples and then generalize them to many data points, drastically reducing the number of annotations.
3.3 Revise
The last and arguably most important step of RioT is integrating the feedback into the model. We apply the general idea of using a loss-based model revision (Schramowski et al., 2020; Ross et al., 2017; Stammer et al., 2020) based on the explanations and the annotation mask. Given the input data $(\mathcal{X},\mathcal{Y})$, we define the original task (or right-answer) loss as $\mathcal{L}_{RA}(\mathcal{X},\mathcal{Y})$. This loss measures the model performance on the classification or forecasting task and is the primary learning objective. To incorporate the feedback, we further use the right-reason loss $\mathcal{L}_{RR}(A,E)$. This loss aligns model explanations $E=\{e(\bm{x})|\bm{x}\in\mathcal{X}\}$ and user feedback $A$. To achieve both model revision and a good task performance, both losses are combined, where $\lambda$ is a hyperparameter to balance both parts of the combined loss.
$\mathcal{L}(\mathcal{X},\mathcal{Y},A,E)=\mathcal{L}_{\mathrm{RA}}(\mathcal{X},\mathcal{Y})+\lambda\mathcal{L}_{\mathrm{RR}}(A,E)$ (3)
Time Domain Feedback. Masking parts of the time domain as feedback is an easy way to mitigate spatially locatable confounders (Fig. 1, top). We use the explanations $E$ and annotations $A$ in the spatial version of the right-reason loss:
$\mathcal{L}^{sp}_{RR}(A,E)=\frac{1}{D}\sum_{\bm{x}\in\mathcal{X}}(e(\bm{x})*a(\bm{x}))^{2}$ (4)
As the explanations and the feedback masks are element-wise multiplied, we effectively train the model to regard the marked parts of the input as unimportant for its output. Thus, using the loss in Eq. 4 as the right-reason component for the combined loss in Eq. 3 allows to effectively steer the model away from points or intervals in time.
Frequency Domain Feedback. However, feedback in the time domain is insufficient to handle every type of confounder. If the confounder is not locatable in time, giving spatial feedback cannot be used to revise the models’ behavior. Therefore, we utilize explanations and feedback in the frequency domain to tackle confounders like in Fig. 1, (bottom). Given the frequency explanations $\hat{E}$ and annotations $\hat{A}$, the right-reason loss for the frequency domain is:
$(\displaystyle\mathcal{L}^{fr}_{RR}(\hat{A},\hat{E})=\frac{1}{D}\sum_{\bm{x}\in\mathcal{X}}\Bigl{(}$$(\text{Re}(\hat{e}(\bm{x}))*\hat{a}_{re}(\bm{x}))^{2}+$ (5)
$(\text{Im}(\hat{e}(\bm{x}))*\hat{a}_{im}(\bm{x}))^{2}\Bigr{)}$
The Fourier transformation is invertible and differentiable, so we can backpropagate directly from this loss. Intuitively, the frequency right-reason loss causes the masked frequency explanations of the model to be small while not affecting any specific point in time.
Depending on the problem at hand, it is possible to use RioT either in the spatial or frequency domain. Moreover, it is also possible to combine feedback in both domains, thus including two right-reason terms in the final loss. This results in two parameters $\lambda_{1}$ and $\lambda_{2}$ to balance the right-answer and both right-reason losses.
$\begin{split}\mathcal{L}(\mathcal{X},\mathcal{Y},A,E)=&\mathcal{L}_{\mathrm{RA}}(\mathcal{X},\mathcal{Y})+\\ &\lambda_{1}\mathcal{L}^{sp}_{\mathrm{RR}}(A,E)+\lambda_{2}\mathcal{L}^{fr}_{\mathrm{RR}}(\hat{A},\hat{E})\end{split}$ (6)
It is important to note that giving feedback in the frequency domain allows a new form of model revision through XIL. Even if we effectively still apply masking in the frequency domain, the effect in the original input domain is entirely different. Masking out a single frequency affects all time points without preventing the model from looking at any of them. In general, having an invertible transformation from the input domain to a different representation allows to give feedback more flexible than before. The Fourier transformation is a prominent example but not the only possible choice for this. Using other transformations like wavelets (Graps, 1995), is also possible.
4 Experimental Evaluations
In this section, we investigate the effectiveness of RioT to mitigate confounders in time series classification and forecasting. Our evaluations include the potential of revising in the spatial domain (RioTsp) and the frequency domain (RioTfreq), as well as both jointly.
4.1 Experimental Setup
Data. We perform experiments on various different datasets. For classification, we focus mainly on the UCR/UEA repository (Dau et al., 2018), which holds a wide variety of datasets for this task. The data originates from different domains, e.g., health records, industrial sensor data, and audio signals. We select all available datasets of a minimal size (cf. App. A.2), which results in Fault Detection A, Ford A, Ford B, and Sleep. We omit experiments on the very small datasets of UCR, as these are generally less suited for deep learning (Ismail Fawaz et al., 2020). For time series forecasting, we evaluate on three popular datasets of the Darts repository (Herzen et al., 2022): ETTM1, Energy and Weather, which all have been used with the TiDE model before (Das et al., 2023). These datasets are sufficiently large, allowing us to investigate the impact of confounding behavior in isolation without the risk of overfitting. We standardize all datasets as suggested by Wu et al. (2021), i.e., rescaling the distribution of values to zero mean and a standard deviation of one.
Production Press Sensor Data (P2S). Our method aims to mitigate real-world confounders in time series data. For a more accurate and reliable assessment of the method’s performance, it is crucial to consider the variations and noise of real-world data and, importantly, real-world confounders. However, there are no available datasets showcasing this problem, challenging models to overcome their confounders. Thus, we introduce the dataset Production Press Sensor Data (P2S) to fill this gap. [^2] https://huggingface.co/datasets/AIML-TUDA/P2S
P2S consists of sensor recordings from a production line for sheet metal parts using a high-speed press. This procedure is one of the sheet metal working industry’s most economically significant manufacturing processes. Specifically, we consider the sensor recordings of the deep bowing phase. Based on the sensor data, the task is to predict a run is normal or has a defect. Experiment runs are conducted with different production speeds, which does not affect the occurrence of defects. However, the speed affects the process’s friction and, thus, the system’s applied forces, leading to confounding factors in the sensor data.
In Fig. 3, this can be clearly seen when comparing the sensor recordings for normal and defect runs at speeds of 80 or 225 strokes per minute. The actual differences between normal and defect runs at the same speed are more subtle; for example, the minimal force can be an indicator. The challenge of the dataset for a model is to pick up on these factors, instead of the deflecting characteristics of the stroke rate. In particular, if no faulty and normal runs of the same speed are in the training data, there is a strong potential for models to focus only on the confounding factors and thus fail to generalize beyond training. Investigating how models can or cannot avoid picking up speed as a confounding factor provides valuable insights into how robust they are to such challenges in the real world. With P2S, we provide a first real-world data basis to evaluate models for their sensitivity to confounding factors in time series data.
Figure 3: Samples of P2S with normal (top) and defect (bottom) setting at 80 and 225 strokes per minute. Areas that vary depending on the stroke rate and are considered confounding and marked red.
The dataset consists of 2264 samples, divided into roughly 200 normal and defect runs of six different speeds. Each of the samples has sensor data of 4096 time steps. In the default setting, the training data has normal and defect runs of different speeds. A reference setting where the model is not affected by any confounder is curated by having normal and defect runs of specifically the same speed in the training set. Further details on the dataset can be found in App. A.4.
Models. For the time series classification datasets, we use the FCN model of Ma et al. (2023). On the dataset Sleep, which was added to UCR after their study, we slightly modified their architecture, which achieves better results (cf. A.1). For forecasting, we use the recently introduced TiDE model (Das et al., 2023). Both models achieve SOTA or at least comparable performance on the respective datasets.
Metrics. In our evaluations, we compare the performance of models on confounded and unconfounded datasets with and without RioT. For classification, we use balanced (multiclass) accuracy (ACC), and for forecasting, we use mean squared error (MSE) and mean absolute error (MAE).
Table 1: Applying RioT mitigates confounders in time series classification. Performance of the FCN model before and after applying RioT for spatial (FCNsp) and frequency (FCNfreq) confounders separately. FCNReference represents the ideal scenario where the model is not affected by any confounder. The high train accuracy and low test accuracy for FCN shows overfitting to the confounder. Applying RioT improves the test performance, indicating its effectiveness in reducing the influence of individual confounders (mean and std over 5 runs).
Config Fault Detection Ford A Ford B Sleep
Train Test Train Test Train Test Train Test
FCNReference 0.99 ±0.00 0.99 ±0.00 0.92 ±0.01 0.91 ±0.00 0.93 ±0.00 0.76 ±0.01 0.68 ±0.00 0.62 ±0.00
FCNsp 1.00 ±0.00 0.74 ±0.06 1.00 ±0.00 0.71 ±0.08 1.00 ±0.00 0.63 ±0.03 1.00 ±0.00 0.10 ±0.03
FCNsp + RioTsp 0.99 ±0.01 0.92 ±0.03 0.99 ±0.01 0.84 ±0.02 0.99 ±0.00 0.68 ±0.02 0.60 ±0.06 0.54 ±0.05
FCNfreq 0.98 ±0.01 0.87 ±0.03 0.98 ±0.00 0.73 ±0.01 0.99 ±0.01 0.60 ±0.01 0.98 ±0.00 0.27 ±0.02
FCNfreq + RioTfreq 0.94 ±0.00 0.90 ±0.03 0.91 ±0.05 0.77 ±0.00 0.97 ±0.00 0.62 ±0.00 0.67 ±0.05 0.44 ±0.06
Table 2: RioT can successfully overcome confounders in time series forecasting. MAE/MSE values on the confounded training set and the unconfounded test set with TiDEReference being the ideal scenario where the model is not affected by any confounder. TiDE shows degraded performance on spatial and frequency confounded datasets, which are mitigated by applying RioT (mean and std over 5 runs).
Metric Config ETTM1 Energy Weather
Train Test Train Test Train Test
$MAE\downarrow$ TiDEReference 0.36 ±0.01 0.48 ±0.01 0.40 ±0.01 0.38 ±0.02 0.16 ±0.01 0.11 ±0.02
TiDEsp 0.32 ±0.01 0.54 ±0.01 0.40 ±0.01 0.85 ±0.01 0.18 ±0.01 0.28 ±0.02
TiDEsp + RioTsp 0.34 ±0.01 0.51 ±0.01 0.57 ±0.01 0.58 ±0.01 0.22 ±0.01 0.18 ±0.00
TiDEfreq 0.18 ±0.01 0.74 ±0.06 0.18 ±0.01 0.53 ±0.07 0.20 ±0.01 0.36 ±0.06
TiDEfreq + RioTfreq 0.19 ±0.01 0.60 ±0.05 0.18 ±0.01 0.40 ±0.03 0.20 ±0.01 0.35 ±0.02
$MSE\downarrow$ TiDEReference 0.27 ±0.01 0.47 ±0.01 0.27 ±0.01 0.26 ±0.02 0.05 ±0.01 0.03 ±0.01
TiDEsp 0.22 ±0.01 0.54 ±0.03 0.28 ±0.01 1.19 ±0.03 0.07 ±0.01 0.14 ±0.02
TiDEsp + RioTsp 0.23 ±0.01 0.48 ±0.01 0.53 ±0.02 0.52 ±0.02 0.10 ±0.01 0.06 ±0.00
TiDEfreq 0.06 ±0.01 0.69 ±0.08 0.07 ±0.01 0.34 ±0.08 0.09 ±0.01 0.16 ±0.04
TiDEfreq + RioTfreq 0.07 ±0.01 0.49 ±0.07 0.07 ±0.01 0.21 ±0.02 0.09 ±0.01 0.15 ±0.02
Confounders. To evaluate how well RioT can mitigate confounders in a more controlled setting, we add shortcuts to the datasets from the UCR and Darts repositories. For classification, we add spatial (FCNsp) or frequency (FCNfreq) shortcut confounders to each sample. The spatial confounder is a distinct pattern per class, replacing the original time series at affected time points. The frequency confounder is a unique frequency per class, which is added to the data, affecting all steps of the time series. The spatial confounder for forecasting (TiDEsp) is also a shortcut confounder, where we copy the forecasting window back into the lookback. This results in a shortcut for every second sample and minor perturbations for the remaining data points. The frequency confounder (TiDEfreq) consists of reoccurring noise in the form of Dirac impulses added to the data. This is not a shortcut confounder but simply distracts from the forecasting task. Confounders are only in the training data, not validation or test data. We generate an annotation mask based on the confounder area or frequency to simulate human feedback. This mask is applied to all confounded samples.
4.2 Evaluations
Removing Confounders for Time Series Classification. First, we evaluate the effectiveness of RioT (RioTsp with spatial and RioTfreq with frequency feedback) in revising confounders in classification tasks. We compare the balanced (multiclass) accuracy of the FCN model in two scenarios: without the application of RioT and with it. As shown in Tab. 1, (top), FCN without RioT overfits to the shortcuts, achieving 100% training accuracy on all datasets but fails to maintain this on the test set. In contrast, applying RioT considerably improves the model’s test performance across all datasets. Although it does not fully reach the upper bound performance of the reference scenario where FCN is trained on unconfounded data, RioT improves performance on average by more than 10%. Additionally, the unconfounded FCN has a notable drop in training-to-test performance on the Ford B dataset, suggesting a shift between training and test distributions. Nevertheless, using RioT enhances the model’s performance on this dataset, too.
Further evaluations of FCN, both with and without RioTfreq, on frequency-confounded data (Tab. 1, bottom) exhibit a pattern similar to that seen with spatial confounders. The model learns the training data shortcuts, achieving near-perfect accuracy, which does not translate to the test set. RioTfreq with frequency feedback again elevates the test performance for all datasets. The least pronounced improvement is on Ford B, with only about a 1% increase when using RioTfreq, leading to the hypothesis that the confounded frequencies in this dataset may carry crucial information, which is inadvertently obscured by the application of RioTfreq.
Removing Confounders for Time Series Forecasting. Confounders are not exclusive to time series classification and can significantly impact other tasks, such as forecasting. In Tab. 2, we demonstrate that confounding factors in the data drastically affect model generalization. The spatially confounded model (TiDEsp) exhibits a notable discrepancy between training and test performance, suggesting it overfitts to the spatial shortcuts. Applying RioTsp to the model lowers both test MAE and MSE on all datasets while having a particularly pronounced effect for Energy of approximately 56% lower test MSE.
In the frequency-confounded setting, the training data includes a recurring noise as a distracting confounder. This again lowers the test performance of TiDE substantially as the model overfits to the training pattern and then fails to generalize. In contrast, using RioTfreq mitigates the model’s excessive focus on the frequency confounder, improving test performances. For instance, in the ETTM1 dataset, the use of RioTfreq decreases the test MAE from 0.54 to 0.51 (a reduction of 5.5%). Similarly, the test MSE decreases from 0.54 to 0.48, representing an 11.1% decrease.
These outcomes from spatial and frequency confounders highlight the efficacy of RioT in addressing different types of confounding factors in time series forecasting. Overall, applying RioT substantially improves test performance for both types of confounders. In some cases, it even reaches the results of the unconfounded reference model, highlighting that RioT fully overcomes the confounder. Interestingly, on the Energy dataset, the performance with RioTfreq even surpasses the reference model. Here, the added frequency has the effect of a form of data augmentation, enhancing the model’s robustness.
Table 3: Applying RioT overcomes the confounder in P2S. Accuracy values on the confounded train set and the unconfounded test set. FCN learns the train confounder, resulting in a drop in test performance. Applying RioT with partial feedback (2) already yields good improvements, while adding feedback on the full confounder area (4) is even better. FCNReference is the ideal scenario where the data is specifically curated so that FCN is not affected by any confounder (mean and std over 5 runs).
P2S Train Test
FCNReference 0.97 ±0.01 0.95 ±0.01
FCN 0.99 ±0.01 0.66 ±0.14
FCN + RioTsp (2) 0.96 ±0.01 0.78 ±0.05
FCN + RioTsp (4) 0.95 ±0.01 0.82 ±0.06
Figure 4: Applying RioT makes the model ignore annotated confounder areas. While FCN primarily focuses on confounder areas, applying RioT with partial feedback (middle) or full feedback (bottom) causes the model to ignore the confounder and focus on the remainder of the input.
Table 4: RioT can combine spatial and frequency feedback. If the data is confounded in the time and frequency domain, RioT can combine feedback on both domains to mitigate confounders. This is superior to applying feedback on only one domain. Reference values represent the ideal scenario when the model is not affected by any confounder (mean and std over 5 runs).
Sleep (Classification ACC $\uparrow$) Train Test
FCNReference 0.68 ±0.00 0.62 ±0.00
FCNfreq,sp 1.00 ±0.00 0.10 ±0.04
FCNfreq,sp + RioTsp 0.94 ±0.00 0.24 ±0.02
FCNfreq,sp + RioTfreq 1.00 ±0.00 0.04 ±0.00
FCNfreq,sp + RioTfreq,sp 0.47 ±0.00 0.48 ±0.03
Energy (Forecasting MSE $\downarrow$) Train Test
TiDEReference 0.28 ±0.01 0.26 ±0.02
TiDEfreq,sp 0.16 ±0.01 0.74 ±0.02
TiDEfreq,sp + RioTsp 0.20 ±0.01 0.61 ±0.02
TiDEfreq,sp + RioTfreq 0.22 ±0.01 0.55 ±0.02
TiDEfreq,sp + RioTfreq,sp 0.25 ±0.01 0.47 ±0.01
Removing Confounders in the Real-World. So far, our experiments have demonstrated the ability to counteract confounders within controlled environments where the confounders were intentionally embedded in the data. However, real-world scenarios often lack this known confounder structure. Our new proposed dataset P2S presents a scenario under real-world conditions. The model explanations for a sample in Fig. 4 (top) reveal a primary focus on four distinct regions of the sensor curve. Using this explanation, one can identify that the model focuses on two regions in the middle. With domain knowledge, it is then possible to state that these regions should not be relevant for predicting whether the outcome is normal or defective. By applying RioT, we can redirect the model’s attention away from these regions. Tab. 3 shows that the initial model is indeed focusing on the confounding run speed, resulting in high training and substantially lower test accuracy. Using RioT to apply feedback to two initially discovered areas already improves test performance substantially (2). However, the explanations of Fig. 4 (middle) reveal that the model still focuses on some confounded input parts. Applying feedback on all four detected regions with RioT improves the test performance even more Tab. 3 (4). While it does not reach the reference setting, where the dataset was specifically curated so that the model does not pick up on the speed confounder in training, applying RioT with two and all four feedback regions shows considerable improvements. Additionally, these results highlight exemplarily the effectiveness of RioT in real-world scenarios where potentially not the full confounder is known (from the beginning).
On P2S, the model has a large variation in test performance, as indicated by the standard deviation in Tab. 3. The variation persists, albeit reduced, even after applying RioT. This suggests intrinsic complexity and variability of how the model learns and adapts to confounders. Nonetheless, applying feedback via RioT consistently narrows the standard deviation, enhancing model reliability and decision-making quality in real-world applications.
Removing Multiple Confounders at Once. In the previous experiments, we illustrated that RioT is suitable for addressing individual confounding factors, whether spatial or frequency-based. Real-world time series data, however, often present a blend of multiple confounding factors that simultaneously may influence model performance. Therefore, we explore the effects of applying RioT simultaneously to spatial and frequency confounders in Tab. 4. When Sleep is confounded both in the spatial and frequency domain, FCN without RioT overfits to the confounders and does not generalize to the test set. Applying only spatial or frequency feedback does not mitigate the confounders, as the model simply learns the other confounder. Combining feedback in both domains (RioTfreq,sp), on the other hand, shows good improvements on the test set, on par with the performance of the previously identified frequency-confounded scenario (cf. Tab. 1). This result aligns with our expectations; given that the original dataset was confounded, the model’s ability to utilize information was limited by the confounders we introduced. These insights not only confirm the utility of RioTfreq,sp in multi-confounder contexts but also suggest a performance upper limit for performance when the underlying data is inherently confounded.
Tab. 4 (bottom) illustrates the impact of multiple confounders on the Energy dataset in the context of forecasting. It is clearly visible that when faced with both spatial shortcut and noise confounders, the model tends to overfit the training data, as indicated by the lower training MSE (corresponding MAE is detailed in the Appendix in Tab. 5). Applying either spatial or frequency feedback individually results in some improvement, as either the spatial shortcut or the frequency noise is avoided. Applying both types of feedback simultaneously (RioTfreq,sp) demonstrates the largest improvements on the test set as feedback on both confounder types is present at the same time.
Nonetheless, the distance between the performance of RioTfreq,sp and the non-confounded model is more pronounced than in the single confounder cases (cf. Tab. 2), suggesting a compounded challenge. This gap indicates the complex nature of interacting confounders, which appears to create a performance upper limit in this case. While RioT is suitable for reducing the confounding effects, the tangled interplay between spatial and frequency confounders may limit the extent to which we can replicate the ideal unconfounded scenario. Research to optimize the deconfounding process in more complex data environments thus represents an important future step.
5 Conclusion
In this work, we present Right on Time a method to mitigate confounding factors in time series data with the help of human feedback. By revising the model, RioT significantly diminishes the influence of these factors, steering the model to align with the correct reasons. Using popular time series models on several manually confounded datasets and the newly introduced, naturally confounded, real-world dataset P2S showcases that they are indeed subject to confounders. Our results, however, demonstrate that applying RioT to these models can mitigate confounders in the data. Furthermore, we have unveiled that addressing solely the time domain is insufficient for revising the model to focus on the correct reasons, which is why we extended our method beyond it. Feedback in the frequency domain provides an additional way to steer the model away from confounding factors and towards the right reasons.
Extending the application of RioT to multivariate time series represents a logical next step, and exploring the integration of various explainer types is another promising direction. Additionally, we aim to apply RioT, especially RioTfreq, to other modalities as well, offering a more nuanced approach to confounder mitigation. It should be noted that while our method shows potential in its current iteration, interpreting attributions in time series data remains a general challenge.
Impact Statement
Our research advances machine learning by enhancing the interpretability and reliability of time series models, significantly impacting human interaction with AI systems. By developing Right on Time (RioT), which guides models to focus on correct reasoning, we improve the transparency and trust in machine learning decisions. While human feedback can provide many benefits, one has also to be aware that it could be incorrect, and evaluate the consequences carefully.
Acknowledgment
This work received funding by the EU project EXPLAIN, funded by the Federal Ministry of Education and Research (grant 01—S22030D). Additionally, this was funded by the project ”The Adaptive Mind” from the Hessian Ministry of Science and the Arts (HMWK), the ”ML2MT” project from the Volkswagen Stiftung, and the Priority Program (SPP) 2422 in the subproject “Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms“ funded by the German Research Foundation (DFG). The latter also contributed the data for the P2S dataset. Furthermore, this work benefited from the HMWK project ”The Third Wave of Artificial Intelligence - 3AI”.
Appendix A Appendix
A.1 Implementation Details
To compute explanations with Integrated Gradients, we follow the common practice of using a baseline of zeros. The standard approach worked well in our experiments, thus we did not explore other baseline choices in this work.
To find suitable parameters for model training, we performed a hyperparameter search over batch size, learning rate and number of training epochs. We then used these parameters for all model trainings and evaluations, with and without RioT. We furthermore selected suitable $\lambda$ values for RioT with a hyperparameter selection on the respective validation sets. The exact values for the model training parameters and the $\lambda$ values can be found in the provided code.
To avoid model overfitting on the forecasting datasets, we performed shifted sampling with a window size of half the lookback window.
A.2 UCR Dataset selection
We focused our evaluation on a subset of UCR datasets with a minimum size. Our selection process was as follows: First, we discarded all multivariate datasets, as we only considered univariate data in this paper. Then we removed all datasets of the category SIMULATED, to avoid datasets which were synthetic from the beginning. We furthermore considered only datasets with less than 10 classes, as having a per-class confounder on more than 10 classes would result in a very high number of different confounders, which would probably rarely happen. Besides these criteria, we discarded all datasets with less than 1000 training samples or a per sample length of less than 100, to avoid the small datasets of UCR. Lastly, we discarded all datasets which could not be loaded, which lead to the resulting four datasets: Fault Detection A, Ford A, Ford B and Sleep.
A.3 Additional Experimental Results
Table 5: RioT can combine spatial and frequency feedback. MAE results when applying feedback in time and frequency with RioT. Combining both feedback domains is superior to feedback on only one of the domains. Reference values represent the ideal scenario when the model is not affected by any confounder (mean and std over 5 runs).
Energy (Forecasting MAE $\downarrow$) Train Test
TiDEReference 0.40 ±0.01 0.38 ±0.02
TiDEfreq,sp 0.30 ±0.01 0.70 ±0.02
TiDEfreq,sp + RioTsp 0.34 ±0.01 0.64 ±0.01
TiDEfreq,sp + RioTfreq 0.36 ±0.01 0.60 ±0.01
TiDEfreq,sp + RioTfreq,sp 0.39 ±0.01 0.55 ±0.01
This section of the appendix provides further insights into our forecasting experiments, specifically showcasing the performance through the MAE metric. The accompanying Tab. 5 details the MAE outcomes for various configurations of our model, emphasizing the effectiveness of integrating spatial and frequency feedback via RioT. These results complement our primary findings (cf. Tab. 4), illustrating the model’s enhanced forecasting performance when addressing confounders across both the time and frequency domains.
A.4 Confounded Dataset from a High-speed Progressive Tool
The presence of confounders is a common challenge in practical settings, affecting models in diverse ways. As the research community strives to identify and mitigate these issues, it becomes imperative to test our methodologies on datasets that mirror the complexities encountered in actual applications. However, for the time domain, datasets with explicitly labeled confounders are not present, highlighting the challenge of assessing model performance against the complex nature of practical confounding factors.
To bridge this gap, we introduce P2S, a dataset that represents a significant step forward by featuring explicitly identified confounders. This dataset originates from experimental work on a production line for deep-drawn sheet metal parts, employing a progressive die on a high-speed press. The sections below detail the experimental approach and the process of data collection.
A.4.1 Real-World setting
The production of parts in multiple progressive forming stages using stamping, deep drawing and bending operations with progressive dies is generally one of the most economically significant manufacturing processes in the sheet metal working industry and enables the production of complex parts on short process routes with consistent quality. For the tests, a four-stage progressive die was used on a Bruderer BSTA 810-145 high-speed press with varied stroke speed. The strip material to be processed is fed into the progressive die by a BSV300 servo feed unit, linked to the cycle of the press, in the stroke movement while the tools are not engaged. The part to be produced remains permanently connected to the sheet strip throughout the entire production run. The stroke height of the tool is 63 mm and the material feed per stroke is 60 mm. The experimental setup with the progressive die set up on the high-speed press is shown in Fig. 5.
Figure 5: Experimental setup with high-speed press and tool as well as trigger for stroke-by-stroke recording of the data
The four stages include a pilot punching stage, a round stamping stage, deep drawing and a cut-out stage. In the first stage, a 3 mm hole is punched in the metal strip. This hole is used by guide pins in the subsequent stages to position the metal strip. During the stroke movement, the pilot pin always engages in the pilot hole first, thus ensuring the positioning accuracy of the components. In the subsequent stage, a circular blank is cut into the sheet metal strip. This is necessary so that the part can be deep-drawn directly from the sheet metal strip. This is a round geometry that forms small arms in the subsequent deep-drawing step that hold the component on the metal strip. In the final stage, the component is then separated from the sheet metal strip and the process cycle is completed. The respective process steps are performed simultaneously so that each stage carries out its respective process with each stroke and therefore a part is produced with each stroke. Fig. 6 shows the upper tool unfolded and the forming stages associated with the respective steps on the continuous sheet metal strip.
Figure 6: Upper tool unfolded and the forming stages associated with the respective steps on the passing sheet metal strip as well as the positions of the piezoelectric force sensors.
A.4.2 Data collection
An indirect piezoelectric force sensor (Kistler 9240A) was integrated into the upper mould mounting plate of the deep-drawing stage for data acquisition. The sensor is located directly above the punch and records not only the indirect process force but also the blank holder forces which are applied by spring assemblies between the upper mounting plate and the blank holder plate. The data is recorded at a sampling frequency of 25 kHz. The material used is DC04 with a width of 50 mm and a thickness of 0.5 mm. The voltage signals from the sensors are digitised using a ”CompactRIO” (NI cRIO 9047) with integrated NI 9215 measuring module (analogue voltage input $\pm$ 10 V). Data recording is started via an inductive proximity switch when the press ram passes below a defined stroke height during the stroke movement and is stopped again as it passes the inductive proximity switch during the return stroke movement. Due to the varying process speed caused by the stroke speeds that have been set, the recorded time series have a different number of data points. Further, there are slight variations in the length of the time series withing one experiment. For this reason, all time series are interpolated to a length of 4096 data points and we discard any time series that deviate considerably from the mean length of a run (i.e., threshold of 3). A total of 12 series of experiments, shown in Tab. 6, were carried out with production rates from 80 to 225 spm. To simulate a defect, the spring hardness of the blank holder was manipulated in the test series that were marked as defect. The manipulated experiments result in the component bursting and tearing during production. In a real production environment, this would lead directly to the parts being rejected.
A.4.3 Data characteristics
Fig. 7 shows the progression of the time series recorded with the indirect force sensor. The force curve characterises the process cycle during a press stroke. The measurement is started by the trigger which is activated by the ram moving downwards. The downholer plates touch down at point A and press the strip material onto the die. Between point A and point B, the downholder springs are compressed, causing the applied force to increase linearly. The deep drawing process begins at point B. At point C, the press reaches its bottom dead centre and the reverse stroke begins so that the punch moves out of the material again. At point D, the deep-drawing punch is released from the material and now the hold-down springs relax linearly up to point E. At point E, the downholder plate lifts off again, the component is fed to the next process step and the process is complete.
Figure 7: Force curve for one stroke. A) set down downholder plate B) start of deep drawing C) bottom dead centre D) deep drawing process completed E) downholder plates lift off F) measurement stops.
Table 6: Overview of conducted runs on the high-speed press with normal and defect states at different stroke rates.
Experiment # State Stroke Rate Samples
1 Normal 80 193
2 Normal 100 193
3 Normal 150 189
4 Normal 175 198
5 Normal 200 194
6 Normal 225 188
7 Defect 80 149
8 Defect 100 193
9 Defect 150 188
10 Defect 175 196
11 Defect 200 193
12 Defect 225 190
Total 2264
A.4.4 Confounders
The presented dataset P2S is confounded by the speed with which the progressive tool is operated. The higher the stroke rate of the press, the more friction is occurring and the higher is the impact of the downholder plate. The differences can be observed in Fig. 3. Since we are aware of these physics-based confounders, we are able to annotate them in our dataset. As the process speed increases, the friction that occurs between the die and the material in the deep-drawing stage changes, as the frictional force is dependent on the frictional speed. This is particularly evident in the present case, as deep-drawing oils, which can optimize the friction condition, were not used in the experiments. The affected area from friction of the punch are in 1380 to 1600 (start of deep drawing) and 2080 to 2500 (end of deep drawing). In addition, the impulse of the downholder plate affecting the die increases due to the increased process dynamics. If the process speed is increased, the process force also increases in the ranges of the time series from 800 to 950 (downholder plate sets down) and 3250 to 3550 (downholder plate lifts).
In the experiment setting of Tab. 3, the training data set is selected in such a way that the stroke rate correlates with the class label, i.e., there are only normal experiments with small stroke rates and defect ones with high stroke rate. Experiment 1, 2, 3, 10, 11, 12 are the training data and the remaining experiments are the test data. To get a reference setting where the model is not affected by any confounder, we use normal and defect experiments with the same speed in training and respectively test data. This results in experiments 1, 3, 5, 7, 9, 11 in the training set and the remaining in the test set.